the length of a table is 11in more than its width. if the area of the table is 1160 square inches, what is its length
Answer by checkley77(12844)

L=W+11
AREA=L*W
1160=(W+11)W
1160=W^2+11W
W^2+11W-1160=0
(W+40)(W-29)=0
W-29=0
W=29 ANS. FOR THE WIDTH.
L=29+11=40 ANS. FOR THE LENGTH.
PROOF:
1160=40*29
1160=1160